# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v3

@@ 4.5 Discussions
4.6 Qualitative Results
In this section, we mainly visualize the qualitative results on RefCOCO val and RefCOCOg val datasets, as shown in Fig. 6 and Fig. 7, respectively. Moreover, we compare the qualitative results of different ways to prompt SAM with texts: (1) our proposed EVF-SAM, (2) SAM with LLM (LISA [20]), and (3) SAM with a CLIP text encoder implemented in this paper (suggested by [1], which are based on the same SAM-Huge model. The qualitative results can demonstrate the superiority of the proposed EVF-SAM.
Visualizations on RefCOCOg. Fig. 7 illustrates the qualitative comparisons on the RefCOCOg val, which aims to segment objects with long expression texts. The SAM with a vanilla CLIP text encoder produces inferior segmentation results given the long-expression texts. However, the proposed EVF-SAM outperforms LISA when using long expressions, even though LISA adopts LLaMA-7B [58] to understand the instructions and generate prompt embeddings, showcasing that the lightweight vision-language models can understand complex expressions. In addition, the proposed EVF-SAM can also understand the texts or expressions towards spatial locations, such as “the umbrella closest to the camera”.
@@ Acknowledgments
This work was partially supported by the National Science and Technology Major Project under Grant No. 2023YFF0905400 and National Natural Science Foundation of China (NSFC) under Grant No. 62276108. We thank Bo Jiang for his helpful feedback on the draft.

### v4

@@ Acknowledgments
This work was partially supported by the National Science and Technology Major Project under Grant No. 2023YFF0905400 and the National Natural Science Foundation of China (NSFC) under Grant No. 62276108. We thank Bo Jiang for his helpful feedback on the draft.
Appendix A Appendix
A.1 Qualitative Results
In this section, we mainly visualize the qualitative results on RefCOCO val and RefCOCOg val datasets, as shown in Fig. 6 and Fig. 7, respectively. Moreover, we compare the qualitative results of different ways to prompt SAM with texts: (1) our proposed EVF-SAM, (2) SAM with LLM (LISA (Lai et al., 2023)), and (3) SAM with a CLIP text encoder implemented in this paper (suggested by (Kirillov et al., 2023), which are based on the same SAM-Huge model. The qualitative results can demonstrate the superiority of the proposed EVF-SAM.
Visualizations on RefCOCOg. Fig. 7 illustrates the qualitative comparisons on the RefCOCOg val, which aims to segment objects with long expression texts. The SAM with a vanilla CLIP text encoder produces inferior segmentation results given the long-expression texts. However, the proposed EVF-SAM outperforms LISA when using long expressions, even though LISA adopts LLaMA-7B (Touvron et al., 2023) to understand the instructions and generate prompt embeddings, showcasing that the lightweight vision-language models can understand complex expressions. In addition, the proposed EVF-SAM can also understand the texts or expressions towards spatial locations, such as ‘the umbrella closest to the camera’.
A.2 Training EVF-SAM with Multi-Tasks
To further enhance the generic capability of our EVF-SAM, we propose to implement multi-task training. Based on the experiments that show the performance degradation when simply including extra segmentation data, we explore ways to make our EVF-SAM gain from extra data.
Mixed training with semantic segmentation. We introduce some extra semantic segmentation datasets (ADE20K (Zhou et al., 2017b), Mapillary (Neuhold et al., 2017)) to proceed with joint training. We do not include COCO-Stuff (Caesar et al., 2018) to avoid data leakage with RefCOCO/+/g. It can be seen in Tab. 7 that the performance on RefCOCO+ and ADE20K gains, indicating the effectiveness of including extra data to enhance the generic capability. However, the evaluation metrics of RefCOCO and RefCOCOg decrease when simply including extra semantic segmentation data. We owe this phenomenon to semantic conflict (Liu et al., 2023e).
Table 7: Results of adding extra semantic data. means zero-shot results. The reported ADE20K results are evaluated on the validation set using the cIoU metric.
ADE20K Mapillary RefCOCO RefCOCO+ RefCOCOg ADE20K
val testA testB val testA testB val test val
82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4   54.2
81.7 83.6 80.3 75.4 78.4 71.3 75.5 77.6 75.9
81.9 83.5 80.3 75.1 78.0 70.8 75.3 77.4   59.6
81.8 83.4 79.7 75.6 78.0 70.7 75.8 76.9 76.1
Unified training with multi-task datasets. To solve the semantic conflict mentioned above, we propose several pre-process strategies for datasets of different distributions. We will open-source related codes in our project page.
- Instance-level data: We apply Objects365 (Shao et al., 2019) to extend RES data. Specifically, (a) for each image, we exclude categories with more than one instance to avoid ambiguity problem. (b) we employ SAM-2 (Ravi et al., 2024) to automatically annotate masks according to the selected ground-truth bounding boxes. The remaining annotations maintain a rich amount thanks to the dense annotation of Objects365 (Shao et al., 2019). We obtain 524K images (of original 600K images) with 1.8M annotations (of original 10M annotations). The mask quality from automatic annotation is fine thanks to the accurate ground-truth from Objects365 (Shao et al., 2019) and the powerful segmentation capability of SAM-2 Ravi et al. (2024). Besides, the remaining annotations are valuable for addressing long-tail problems because those excluded annotations often belong to head categories.
- Semantic-level data: We introduce ADE20K (Zhou et al., 2017a; 2019) to broaden multi-task capability. We construct a special token ‘[semantic]’ and input ‘[semantic] {category}’. The special token would not be limited to common grammar so it is helpful to avoid semantic conflict.
- Part-level data: To enable the model to segment parts of objects, we introduce PartImageNet (He et al., 2022), HumanParsing (Liang et al., 2015a; b) and PASCAL-Part (Chen et al., 2014) to train our model. For semantic-level annotated datasets, i.e., HumanParsing, we implement the same strategy as ADE20K. Exceptionally, we align the definition of ‘left’ and ‘right’ with RES datasets (e.g., RefCOCO). For instance-level annotated datasets, i.e., PartImageNet and PASCAL-Part, we merge instance masks of the same category to convert the dataset to semantic-level. Then, the same strategy as ADE20K is implemented.
By combining those datasets, we observe a significant performance gain of 1.0 cIoU on the average metric, as shown in Tab. 2. Moreover, our model is able to proceed with multiple tasks like part-segmentation and semantic-level segmentation.